OpenStudy (anonymous):

determine the vertex and the axis of symmetry for the fraction below y=-x^2+2x+1

OpenStudy (anonymous):

vertex of a parabola \(y=ax^2+bx+c\) always has the first coordinate \(-\frac{b}{2a}\) which in your case is \(-\frac{2}{2\times (-1)}=1\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by \(1\)

OpenStudy (anonymous):

So the vertex s is -1 2 and the symmetry is x=1

OpenStudy (anonymous):

no the first coordinate of the vertex is \(1\) not \(-1\)

OpenStudy (anonymous):

the axis of symmetry is \(x=1\) though, you are right for that one

OpenStudy (anonymous):

and the second coordinate is 2 as you said

OpenStudy (stamp):

see attachment

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

it is \((1,2)\) but all else is right yw